Beplay stock is selling for 44.38? Their long run growth rate is 6.8% and their required return return is 12.6%. What was the most recent dividend paid by Beplay?
Answer #1

HI,

Here stock price P = 44.38

growth rate g= 6.8%

required return r = 12.6%

most recent dividend paid D0=?

as per dividend discount model

stock price P= D0*(1+g)/(r-g)

44.38 = D0*(1+6.8%)/(12.6%-6.8%)

44.38*0.058=D0*1.068

2.574 = D0*1.068

D0 = $2.41

Hence most recent dividend is $2.41

Thanks